In mid-December 2001, Jim Vala has another meeting with Bill Jones.

During the meeting, Jones says:

“As you know, Jones Group has recently divested DCom. At the time of that transaction,

we signed a 20-year lease that gives DCom access to our pipeline property. We have just

been informed that, as a result of current market conditions, DCom will default on the

lease payments. We need that cash to service the debt that was retained by Jones Group

at the time of the divestiture.”

Jones continues:

“We should consider a new dividend policy to conserve cash. Dividends per share will

total $1.85 in 2001. I think we should reduce our dividend by 60 percent and maintain

that lower level for 2002 and 2003 to allow us to pay off some debt. In 2004, we will

increase our dividend back to $1.85, then grow the dividend at 8 percent annually

thereafter.”

The required rate of return on Jones Group equity is 11 percent for the foreseeable future.

Calculate, using a dividend discount model (DDM) approach, the expected share price of Jones

Group equity on January 1, 2002, given the new dividend policy described by Jones. Show your

calculations.
Answer #1

Hello,

In this case we will use multi period dividend approach. Please see below snapshot for solutionParticulars Dividend Dividend Growth Rate Perpetual Dividend Perpetual Growth Rate Required Rate of Return Total Dividends Di

DEF =C3+1 =C4*(1+D5) -0.6 =D3+1 =D4*(1+E5) 1.85 =E3+1 11.85 =F4/E4-1 =(F4*(1+07))/(CE Particulars Dividend Dividend Growth Ra